# Randomized methods for dynamical low-rank approximation

### v2

@@ 2.3 Adaptive dynamical rangefinder
In most applications, the rank is not known a priori. Instead, one might prefer to prescribe an error tolerance and let the method find an appropriate rank such that the error $\left\|A(h)-Q_{h}Q_{h}^{T}A(h)\right\|_{2}$ will be smaller than the tolerance. The so-called adaptive randomized rangefinder [HMT11, Algorithm 4.2] relies on a lemma from [WLRT08]:
@@
$A(h)\omega^{(i)}=A(h)\Omega_{[:,i]}\approx B(h)_{[:,i]}.$
In other words, $B(h)$ can be used as an accurate error estimate at little extra cost. We define the adaptive dynamical rangefinder algorithm as follows:
@@
    $&\accentset{\mbox{\large.}}{B}(t)=F(B(t)\Omega^{\dagger})\Omega,\quad t\in[0,h]\\ &B(0)=A_{0}\Omega.\end{aligned}\right.$
@@
    Compute the error estimate $E=\max_{i=1,\ldots,K}\left\|B(h)_{[:,i]}-Q_{h}Q_{h}^{T}B(h)_{[:,i]}\right\|_{2}$.
@@
$E=\max_{i=1,\ldots,K}\left\|B(h)_{[:,i]}-Q_{h}Q_{h}^{T}B(h)_{[:,i]}\right\|_{2}.$
@@
12:         Set $j=j+1$ and update the basis: $Q_{h}=\mathrm{orth}([Q_{h},B(h)])$triangleright e.g. by a QR update
@@ 2.4 Finding the range: a toy experiment
and can therefore be used to compare the idealized and practical methods above defined. In the following, we set the parameters to $n=100$ and $h=0.1$. On one hand, we perform the idealized method by applying the rangefinder algorithm to the closed form solution $X(h)$ and we extract the matrix $Q$ with orthogonal columns. On the other hand, we perform the new dynamical rangefinder algorithm applied to equation (10) and we extract the matrix $Q_{h}$ with orthogonal columns. We emphasize that in the dynamical rangefinder algorithm, we never solve the full order model but only the reduced differential equation as defined in (7). The target rank is $r=5$ and the oversampling parameter $p$ varies from 00 to $12$. Due to the random nature of the algorithm, we perform $100$ simulations and show a box-plot showing the relative errors made by each method compared to the reference solution
@@
Figure 1: Application to the toy problem (10). The target rank of approximation is $r=5$ and the figure shows the relative error (at time $h=0.1$) made by the methods as we increase the oversampling parameter. The boxes show the median, first and third quartiles and extreme values of the $100$ simulations.
@@ 3 Time-stepping methods
In the following, we will assume that the problem can be well-approximated by a low-rank approximation, and in particular we will start from an approximate of the initial value $A_{0}\approx Y_{0}=U_{0}S_{0}V_{0}^{T}\in\mathcal{M}_{r}$ where $U_{0}\in\mathbb{R}^{m\times r}$ and $V_{0}\in\mathbb{R}^{n\times r}$ have orthogonal columns, and $S_{0}\in\mathbb{R}^{r\times r}$ is invertible. We could take for example $Y_{0}=\mathcal{T}_{r}(A_{0})$, or an approximate of it. We fix a stepsize $h>0$ and a target rank of approximation $r$, the goal is to find factors such that $A(h)\approx Y_{1}=U_{1}S_{1}V_{1}^{T}\in\mathcal{M}_{r}$ where $U_{1}$ and $V_{1}$ have orthogonal columns. Building on the newly defined dynamical rangefinder algorithm, we propose two new time-stepping methods which evolve such a low-rank approximation and which do not involve the projected dynamics (4). The new time-stepping methods are defined by one step, which has to be repeated to get an approximate solution on the whole time interval. Ultimately, we discuss a simple modification of the two methods that allows to update and automatically choose the rank at each time step.
@@ 3.1 Dynamical randomized SVD
Let us start by motivating the method with an idealized procedure. Consider a target rank $r$ and an oversampling parameter $p$ and let $\mathcal{Q}_{h}\in\mathbb{R}^{m\times(r+p)}$ be the matrix with orthogonal columns obtained from the idealized dynamical rangefinder method described in Section 2.1. We propose the following idealized procedure:
@@
    $&\accentset{\mbox{\large.}}{\mathcal{C}}(t)=F(A(t))^{T}\mathcal{Q}_{h},\quad t\in[0,h]\\ &\mathcal{C}(0)=A_{0}^{T}\mathcal{Q}_{h}.\end{aligned}\right.$ (12)
@@
Solving the differential equation (12) requires access to the solution to the full order model, and is therefore not feasible in practice. Instead, we propose to approximate the solution by projecting the differential equation on an approximate range.
@@
Because of the time discretization, approximating equation (12) on the estimated range at time $h$ might lead to a poor numerical approximation. We therefore need to augment the approximation space with the information from the initial value. Doing so increases the computational cost of the method but leads to better numerical results. Starting from the initial approximation $A_{0}\approx Y_{0}=U_{0}S_{0}V_{0}^{T}\in\mathcal{M}_{r}$, we define one step of the dynamical randomized SVD (DRSVD) as follows:
@@
    $&\accentset{\mbox{\large.}}{C}(t)=F(QC(t)^{T})^{T}Q,\quad t\in[0,h]\\ &C(0)=Y_{0}^{T}Q.\end{aligned}\right.$ (14)
@@
4:Compute $Q_{h}=\mathrm{DynamicalRangefinder}(F,Y_{0},h,r,p,q)$triangleright See Algorithm 1
5:Augment the basis: $Q=\mathrm{orth}([U_{0},Q_{h}])$triangleright With pivoted QR, or else
@@ 3.2 Dynamical generalized Nyström
Similarly to the dynamical randomized SVD presented above, we could give an idealized generalized Nyström algorithm based on the exact solution, which would also verify the exactness property stated in Lemma 3.1. The ideas being very similar, we skip the discussion of the idealized method and present directly the practical algorithm. Consider a target rank $r>0$, two oversampling parameters $p$, $\ell\geq 0$, and a low-rank approximate of the initial value $A_{0}\approx Y_{0}=U_{0}S_{0}V_{0}^{T}\in\mathcal{M}_{r}$. We define one step of the dynamical generalized Nyström (DGN) algorithm as follows:
@@
    $&\accentset{\mbox{\large.}}{B}(t)=F(B(t)W^{T})W,\\ &B(0)=A_{0}W,\end{aligned}\right.\qquad\left\{\begin{aligned} &\accentset{\mbox{\large.}}{C}(t)=F(QC(t)^{T})^{T}Q,\\ &C(0)=A_{0}^{T}Q,\end{aligned}\right.\qquad\left\{\begin{aligned} &\accentset{\mbox{\large.}}{D}(t)=Q^{T}F(QD(t)W^{T})W,\\ &D(0)=Q^{T}A_{0}W.\end{aligned}\right.$ (16)
@@
$\accentset{\mbox{\large.}}{B}(t)=F(B(t)W^{T})W,\quad B(0)=Y_{0}W.$
@@
$\accentset{\mbox{\large.}}{C}(t)=F(QC(t)^{T})^{T}Q,\quad C(0)=Y_{0}^{T}Q.$
@@
$\accentset{\mbox{\large.}}{D}(t)=Q^{T}F(QD(t)W^{T})W,\quad D(0)=Q^{T}Y_{0}W.$
@@ 3.3 Rank-adaptive time-stepping methods
In the two new methods above presented, a fixed target rank is prescribed, which assumes that we have a guess on the numerical rank of the solution. In practice, we do not have always such a guess and we typically need to change the rank at each time step. Unlike the projector-splitting integrator, the augmented space together with the oversampling parameter $p$ already allow some freedom in the truncation step so that the rank can change from one step to the next. However, the rank growth is then limited to $2r+p$, which might not be enough to catch fast changing solutions. Instead, we propose a more systematic approach based on the adaptive dynamical rangefinder algorithm proposed in Section 2.3, which allows to choose automatically the rank.
Consider a prescribed error tolerance $\tau>0$ and a certain probability of failure $\beta>0$. Let us denote the tolerance-based truncated SVD by $\mathcal{T}_{\tau}$ such that
@@ 4.1 A stiff differential equation with Lyapunov structure
with Dirichlet boundary conditions and where $(x,y)\in\Omega$ and $t\in[0,T]$. Here, the source $s(x,y)$ is constant in time. A spatial discretization leads to the Lyapunov differential equation [UV20] defined by
@@
where $A(t)\in\mathbb{R}^{n\times n}$, $\alpha\geq 0$ and $t\in[0,T]$. For our numerical experiments, we use $\Omega=[-\pi,\pi]^{2}$ discretized with $n=256$ points. The discrete Laplacian operator is $L=(n-1)^{2}\cdot\mathrm{diag}(1,-2,1)\in\mathbb{R}^{n\times n}$. The construction of the initial value and the source is similar to [UV20, LCK24]. Since Assumption 2 is not satisfied at the initial value, we propagate the solution from time 00 to $10^{-4}$ and start our numerical experiments from there. We point out that Assumption 2 is only required for DLRA techniques, and we do not require it for the new methods. The new methods do not require this extra propagation of the initial value.
In Table 1, we compare the accuracy of the methods applied to equation (21) with a target rank $r=5$. When the stepsize is large (here $h=0.1$), existing DLRA methods perform poorly because the problem is stiff. The new methods perform well in comparison with a good accuracy and low variance despite the randomness. Note that the DGN algorithm with $q=1$ power iteration is able to recover the reference best rank approximation.
In Figure 2, we compare the time of computation of the new methods with the augmented BUG integrator. We can see here the large hidden constant of the augmented BUG integrator: with a stepsize of $h=0.1$ the method has not converged. We therefore need a smaller stepsize, here $h=0.01$, to get convergence of the augmented BUG. We point out that, for the same accuracy, the new methods are faster than the augmented BUG and the DGN can reach a higher accuracy when the rank of approximation is large enough.
@@
Figure 2: Accuracy and computation time (in seconds) of the methods applied to the Lyapunov equation (21). The integers next to the markers correspond to the target rank. The reference solver is a second-order exponential Runge–Kutta method with stepsize $\delta t=10^{-4}$, and the substeps were solved with the reference solver.
@@ 4.2 Reaction-diffusion with the Allen–Cahn equation
the domain is $\Omega=[0,2\pi]^{2}$ and the spatial mesh contains $128\times 128$ points with periodic boundary conditions. After finite differences, we obtain the matrix differential equation
@@
where the power is taken element-wise. We consider the time interval $[0,T]=[0,10]$, and most of the reaction occurs between $t=3$ and $t=7$. At the end of the time interval, the solution has almost reached the steady-state. At any given time, the singular values of the solution decay exponentially fast.
@@ 4.3 Stochastic problems with the Burgers equation
$\partial_{t}v+v\partial_{x}v=\nu\partial_{xx}v,\quad x\in\Omega,\quad t\in[0,T],$ (24)
@@
The problem is discretized by associating each column to a random sample. Hence, taking the space $\Omega=[0,1]$ and using central finite differences leads to the matrix differential equation
@@
where the unknown matrix is $A(t)\in\mathbb{R}^{n\times s}$ and the operation * is the element-wise (Hadamard) product. The discrete differential operators are
@@
Figure 4: Relative error over time of several methods applied to the stochastic Burgers equation (26) with stepsize $h=0.01$ and $30$ simulations. The reference solver is scipy RK45 method with absolute and relative tolerance set to $10^{-12}$, and the substeps are solved with the reference solver.

### v3

@@ 2.3 Adaptive dynamical rangefinder
In most applications, the rank is not known a priori. In a rank adaptive method, we prescribe an error tolerance and let the method choose a suitable rank such that the error $\left\|A(h)-Q_{h}Q_{h}^{T}A(h)\right\|_{2}$ will be smaller than the desired tolerance. The so-called adaptive randomized rangefinder [HMT11, Algorithm 4.2] relies on a lemma from [WLRT08]:
@@
$A(h)\omega^{(i)}=A(h)\Omega[:,i]\approx B(h)[:,i]\qquad\text{ where }[:,i]\text{ denotes the $i$th column.}$
In other words, the columns of $B(h)$ can be used as an accurate error estimate at little extra cost. We define the adaptive dynamical rangefinder algorithm as follows:
@@
    $\left\{\begin{aligned} &\accentset{\mbox{\large.}}{B}(t)=F(B(t)\Omega^{\dagger})\Omega,\quad t\in[0,h]\\ &B(0)=A_{0}\Omega.\end{aligned}\right$.
@@
    Compute the error estimate $E=\max_{i=1,\ldots,K}\left\|B(h)[:,i]-Q_{h}Q_{h}^{T}B(h)[:,i]\right\|_{2}$.
@@
$E=\max_{i=1,\ldots,K}\left\|B(h)[:,i]-Q_{h}Q_{h}^{T}B(h)[:,i]\right\|_{2}.$
@@
12:   Set $j=j+1$ and update the basis: $Q_{h}=\mathrm{orth}([Q_{h},B(h)])$ $\triangleright$ e.g. by a QR update
@@ 2.4 Finding the range: a toy experiment
which we use to compare the idealized and practical methods above defined. In the following, we set the parameters to $n=100$ and $h=0.1$. On the one hand, we perform the idealized method by applying the rangefinder algorithm to the closed form solution $X(h)$ and we extract the matrix $Q$ with orthonormal columns with a QR decomposition. On the other hand, we perform the new dynamical rangefinder algorithm 1 applied to equation (10) and call $Q_{h}$ the resulting matrix with orthonormal columns. We emphasize that with the dynamical rangefinder algorithm, we never solve the full order model but only the reduced differential equation as defined in (7). The target rank is $r=5$ and the oversampling parameter $p$ varies from 0 to $12$. Due to the random nature of the algorithm, we perform $100$ simulations and show a box-plot showing the relative errors made by each method compared to the reference solution
@@
Figure 1: Comparison of the dynamical rangefinder algorithm 1 with the rangefinder [HMT11] applied to the toy problem 10. The target rank of approximation is $r=5$ and the figure shows the relative error (at time $h=0.1$) made by the methods as we increase the oversampling parameter. The boxes show the median, first and third quartiles and extreme values of the $100$ simulations.
@@ 3 Time-stepping methods
In the following, we will assume that the problem can be well-approximated by a low-rank approximation, and in particular we will start from an approximate of the initial value $A_{0}\approx Y_{0}=U_{0}S_{0}V_{0}^{T}\in\mathcal{M}_{r}$ where $U_{0}\in\mathbb{R}^{m\times r}$ and $V_{0}\in\mathbb{R}^{n\times r}$ have orthonormal columns, and $S_{0}\in\mathbb{R}^{r\times r}$ is invertible. We could take for example $Y_{0}=\mathcal{T}_{r}(A_{0})$, or an approximate of it. We fix a stepsize $h>0$ and a target rank of approximation $r$, the goal is to find factors such that $A(h)\approx Y_{1}=U_{1}S_{1}V_{1}^{T}\in\mathcal{M}_{r}$ where $U_{1}$ and $V_{1}$ have orthonormal columns. Building on top of the new dynamical rangefinder algorithm, we propose two new time-stepping methods which evolve such a low-rank approximation and which do not involve the projected dynamics (4). We only present one step of each method, which the time-stepping method will repeat to get an approximate solution on the whole time interval. Ultimately, we discuss a simple modification of the two methods that allows to update and automatically choose the rank at each time step.
@@ 3.1 Dynamical randomized SVD
Let us start by motivating the method with an idealized procedure. Consider a target rank $r$ and an oversampling parameter $p$ and let $\mathcal{Q}_{h}\in\mathbb{R}^{m\times(r+p)}$ be the matrix with orthonormal columns obtained from the idealized dynamical rangefinder method described in Section 2.1. We propose the following idealized procedure:
@@
    $\left\{\begin{aligned} &\accentset{\mbox{\large.}}{\mathcal{C}}(t)=F(A(t))^{T}\mathcal{Q}_{h},\quad t\in[0,h]\\ &\mathcal{C}(0)=A_{0}^{T}\mathcal{Q}_{h}.\end{aligned}\right$. (12)
@@
Solving differential equation (12) requires access to the full order model solution, and is therefore not feasible in practice. Instead, we propose to approximate the solution by projecting the differential equation on an approximate range.
@@
Because of the time discretization, approximating equation (12) on the estimated range at time $h$ might lead to a poor numerical approximation. We therefore need to augment the approximation space with the information from the initial value. Doing so increases the computational cost of the method but also increase its accuracy. Starting from the initial approximation $A_{0}\approx Y_{0}=U_{0}S_{0}V_{0}^{T}\in\mathcal{M}_{r}$, we define one step of the dynamical randomized SVD (DRSVD) as follows:
@@
    $\left\{\begin{aligned} &\accentset{\mbox{\large.}}{C}(t)=F(QC(t)^{T})^{T}Q,\quad t\in[0,h]\\ &C(0)=Y_{0}^{T}Q.\end{aligned}\right$. (14)
@@
4:Compute $Q_{h}=\mathrm{DynamicalRangefinder}(F,Y_{0},h,r,p,q)$ $\triangleright$ See Algorithm 1
5:Augment the basis: $Q=\mathrm{orth}([U_{0},Q_{h}])$ $\triangleright$ With pivoted QR, or else
@@ 3.2 Dynamical generalized Nyström
Similarly to the dynamical randomized SVD presented above, we could give an idealized generalized Nyström algorithm based on the exact solution, which would also verify the exactness property stated in Lemma 3.1. The ideas being very similar, we skip the discussion of the idealized method and present only the practical algorithm. Consider a target rank $r>0$, two oversampling parameters $p$, $\ell\geq 0$, and a low-rank approximate of the initial value $A_{0}\approx Y_{0}=U_{0}S_{0}V_{0}^{T}\in\mathcal{M}_{r}$. We define one step of the dynamical generalized Nyström (DGN) algorithm as follows:
@@
    $\left\{\begin{aligned} &\accentset{\mbox{\large.}}{B}(t)=F(B(t)W^{T})W,\\ &B(0)=A_{0}W,\end{aligned}\right.\qquad\left\{\begin{aligned} &\accentset{\mbox{\large.}}{C}(t)=F(QC(t)^{T})^{T}Q,\\ &C(0)=A_{0}^{T}Q,\end{aligned}\right.\qquad\left\{\begin{aligned} &\accentset{\mbox{\large.}}{D}(t)=Q^{T}F(QD(t)W^{T})W,\\ &D(0)=Q^{T}A_{0}W.\end{aligned}\right$. (16)
@@
$\accentset{\mbox{\large.}}{B}(t)=F(B(t)W^{T})W,\quad B(0)=Y_{0}W,\qquad t\in[0,h].$
@@
$\accentset{\mbox{\large.}}{C}(t)=F(QC(t)^{T})^{T}Q,\quad C(0)=Y_{0}^{T}Q,\qquad t\in[0,h].$
@@
$\accentset{\mbox{\large.}}{D}(t)=Q^{T}F(QD(t)W^{T})W,\quad D(0)=Q^{T}Y_{0}W,\qquad t\in[0,h].$
@@ 3.3 Rank-adaptive time-stepping methods
In the two new methods above presented, a fixed target rank is prescribed, which assumes that we have a guess on the numerical rank of the solution. In practice, we do not have always such a guess and we typically need to change the rank at each time step. Unlike the projector-splitting integrator, the augmented space together with the oversampling parameter $p$ already allow some freedom in the truncation step so that the rank could evolve from one step to the next. However, the rank growth would be limited to $2r+p$, which might not be enough to catch fast changing solutions. Instead, we propose a more systematic approach based on the adaptive dynamical rangefinder algorithm proposed in Section 2.3, which allows to choose automatically the rank.
Consider a prescribed error tolerance $\tau>0$ and a certain probability of failure $\beta>0$. Let us denote the tolerance-based truncated SVD by $\mathcal{T}_{\tau}$ and defined such that
@@ 4.1 A stiff differential equation with Lyapunov structure
with Dirichlet boundary conditions and where $(x,y)\in\mathcal{D}$ and $t\in[0,T]$. Here, the source $s(x,y)$ is constant in time. A spatial discretization via finite differences leads to the Lyapunov differential equation [UV20] defined by
@@
where $A(t)\in\mathbb{R}^{n\times n}$, $\alpha\geq 0$ and $t\in[0,T]$. For our numerical experiments, we use $\Omega=[-\pi,\pi]^{2}$ discretized with $n=256$ points. The discrete Laplacian operator is $L=(n-1)^{2}\cdot\mathrm{diag}(1,-2,1)\in\mathbb{R}^{n\times n}$. The construction of the initial value and the source is similar to [UV20, LCK24]. Since Assumption 2 is not satisfied at the initial value, we propagate the solution from time 0 to $10^{-4}$ and start our numerical experiments from there. We point out that Assumption 2 is only required for DLRA techniques, and we do not require it for the new methods. The new methods do not require this extra propagation of the initial value.
In Table 1, we compare the accuracy of the methods applied to equation (21) with a target rank $r=5$. When the stepsize is large (here $h=0.1$), existing DLRA methods perform poorly because the problem is stiff. The new methods perform well in comparison with a good accuracy and low variance despite the randomness. The table shows the importance of the power iteration for the DRSVD algorithm. Note that the DGN algorithm with $q=1$ power iteration is able to recover the reference best rank approximation.
In Figure 2, we report the computational time versus accuracy for the augmented BUG integrator with two different step sizes and for the proposed methods. For $h=0.1$, the augmented BUG is fastest but fails to converge, as the relative error stagnates above our expectations (for example by comparing the errors obtained with a target rank $r=2$). Reducing the stepsize to $h=0.01$ restores convergence, but at the expense of substantially increased computational cost. By contrast, the proposed methods (DRSVD and DGN) achieve significantly higher accuracy at lower cost thanks to a convergence achieved with a larger stepsize. In particular, DGN consistently reaches the highest attainable accuracy (down to $10^{-15}$) when the target rank is sufficiently large, thereby outperforming the augmented BUG integrator in both efficiency and robustness to get convergence.
@@
Figure 2: Accuracy and computation time (in seconds) of the methods applied to the Lyapunov equation (21). The integers next to the markers correspond to the target rank. The reference solver is a second-order exponential Runge–Kutta method with stepsize $\delta t=10^{-4}$, and the substeps of the low-rank methods were solved with this reference solver.
@@ 4.2 Reaction-diffusion with the Allen–Cahn equation
the domain is $\mathcal{D}=[0,2\pi]^{2}$ and the spatial mesh contains $128\times 128$ points with periodic boundary conditions. After finite differences, we obtain the matrix differential equation
@@
where the power is taken element-wise (Hadamard product). We consider the time interval $[0,T]=[0,10]$, and most of the reaction occurs between $t=3$ and $t=7$. At the end of the time interval, the solution has almost reached the steady-state. At any given time, the singular values of the solution decay exponentially fast.
@@ 4.3 Stochastic problems with the Burgers equation
$\partial_{t}v+v\partial_{x}v=\nu\partial_{xx}v,\quad x\in\mathcal{D},\quad t\in[0,T],$ (24)
@@
The problem is discretized by associating each column to a random sample. Hence, taking the space $\mathcal{D}=[0,1]$ and using central finite differences leads to the matrix differential equation
@@
where the unknown matrix is $A(t)\in\mathbb{R}^{n\times s}$ and the operation * is again the element-wise (Hadamard) product. The discrete differential operators are
@@
Figure 4: Relative error over time of several methods applied to the stochastic Burgers equation (26) with stepsize $h=0.01$ and $30$ simulations. The reference solver is scipy RK45 method with absolute and relative tolerance set to $10^{-12}$, and the substeps of the low-rank methods are solved with this reference solver.